# Liquid: Language Models are Scalable Multi-modal Generators

### v1

@@ Abstract
We present Liquid, an auto-regressive generation paradigm that seamlessly integrates visual comprehension and generation by tokenizing images into discrete codes and learning these code embeddings alongside text tokens within a shared feature space for both vision and language. Unlike previous multimodal large language model (MLLM), Liquid achieves this integration using a single large language model (LLM), eliminating the need for external pretrained visual embeddings such as CLIP. For the first time, Liquid uncovers a scaling law that performance drop unavoidably brought by the unified training of visual and language tasks diminishes as the model size increases. Furthermore, the unified token space enables visual generation and comprehension tasks to mutually enhance each other, effectively removing the typical interference seen in earlier models. We show that existing LLMs can serve as strong foundations for Liquid, saving 100× in training costs while outperforming Chameleon in multimodal capabilities and maintaining language performance comparable to mainstream LLMs like LLAMA2. Liquid also outperforms models like SD v2.1 and SD-XL (FID of 5.47 on MJHQ-30K), excelling in both vision-language and text-only tasks. This work demonstrates that LLMs such as LLAMA3.2 and GEMMA2 are powerful multimodal generators, offering a scalable solution for enhancing both vision-language understanding and generation. The code and models will be released.
@@ 1 Introduction
(a) MMLU zero-shot performance
(b) Mean score of language tasks
(c) VQAScore in GenAI-Bench
Figure 1: The performance impact of multi-modal mixed training versus text-only training on language tasks. Figure (a) compares the performance of both methods on MMLU (zero-shot), while Figure (b) shows the average scores on 5 language tasks, including HellaSwag, WinoGrande, ARC-Easy, ARC-Challenge, and BoolQ. It can be seen that multi-modal mixed training does impact language performance when the model size is small. However, this degradation gradually disappears as the model size increases, demonstrating that larger LLMs have greater capacity to learn visual generation without impacting the model’s original capabilities. Figure (c) demonstrates that the image generation capabilities of the model are continuously improved as the model size increases. We report the VQA score on the advanced prompts set of GenAI-Bench.
@@
In light of the aforementioned issues, an emerging line of research [15, 49, 12, 71, 62, 56] attempts to employ the VQVAE [65, 15] model as a universal visual tokenizer for MLLMs. Analogous to the role played by the BPE [53] tokenizer in LLMs, VQVAE establishes a bi-directional mapping between raw pixels and discrete codes. This enables the MLLMs to learn visual code embeddings jointly with text tokens, rather than constrained by the feature space of a pretrained visual encoders like CLIP or pretrained visual generators like diffusion models. Moreover, the discrete nature of visual tokens allows uniform modeling of visual and text tokens with the same next-token prediction loss, which seamlessly integrates both modalities. LWM [39] and Chameleon [59] are the pioneers in exploring this approach. However, these methods involve extensive training from scratch, which makes it computationally expensive to disseminate exploration in this form. Follow-up works further introduces bi-directional attention [75], discrete diffusion modeling [69], and residual codes [68] to handle visual tokens. Although these techniques improve the performance in visual generation and understanding, they unavoidably impose discrepancies in vision and language modeling.
In this paper, we revisit the plain design of MLLMs and present Liquid, a scalable decoder-only architecture for multi-modal generation and understanding. We employ the VQVAE [15] model as the visual tokenizer to encode images into discrete codes, which share the same vocabulary and embedding space with text tokens. Such a strategy ensures that Liquid can benefit from the mature structure designs and acceleration techniques of LLMs. We find that existing LLMs are excellent starting points for training, as they have already acquired strong semantic understanding and generation capabilities. We directly build our model on existing LLMs [60, 61, 14, 24] and stick with the original structure of LLMs, solely expanding their vocabulary size to accommodate the image token indices derived from VQVAE. We employ both text-only and image-text pair data to conduct multi-modal mixed training on Liquid, thereby enabling it to simultaneously continue functioning on text-only tasks, image understanding tasks, and text-guided image generation tasks. Compared to Chameleon [59], which is trained from scratch on massive amounts of low-quality data, we continue training existing LLMs on a smaller amount of high-quality data to equip them with visual understanding and generation capabilities. This approach saves 100× of training cost while achieving stronger multimodal capabilities and maintains language capabilities comparable to mainstream LLMs.
We further explore the scaling property and performance of Liquid with different scales of LLMs. We find that when visual and language tasks are trained in a unified space, smaller models experience a performance drop in the original language tasks and perform worse in visual generation tasks compared to models trained solely on visual generation. However, as the model size increases, the performance trade-off gradually disappears. This phenomenon demonstrates that larger LLMs have sufficient capacity to handle both visual generation and language tasks simultaneously. This highlights the scaling advantage of utilizing LLMs directly for image generation. To our knowledge, no previous works have studied the impact of unifying these tasks in an auto-regressive form. Notably, through data ablation experiments, we discover that the training of visual understanding and visual generation can mutually boost each other under the framework of the Liquid. It further confirms the tremendous potential of unifying visual generation and understanding tasks into LLMs.
@@
  • Comprehensive experiments about scaling laws of unified multi-modal models, revealing that the trade-off between language and visual tasks diminishes as scale of model increases, thereby proving the applicability of the appealing properties of LLMs.
@@ 2 Related Work
3 Method
3.1 Image Tokenizer
We use the same tokenizer as Chameleon [59], which is a VQGAN [15]. It encodes a 512 × 512 image into 1024 discrete tokens in a codebook of size 8192. We appended these discrete image tokens to the text codebook produced by the BPE [53] tokenizer to expand the vocabulary size of the LLM, thereby extending its language space to a multi-modal space encompassing both visual and linguistic elements. To speed up training, we use this tokenizer to pre-convert all images used during training into discrete tokens.
3.2 Architecture
As illustrated in Fig. 3, Liquid can be constructed based on any existing large language models (LLMs). In this paper, we employ the GEMMA-7B [60] and GEMMA-2 series [61] as the base model of Liquid. We refrain from altering any structures within the LLMs to facilitate continued training directly from pre-trained weights. The only modification is the addition of 8192 new learnable embeddings for discrete image tokens. Correspondingly, we extend the original LM head by 8192 dimensions to enable the model to predict both text and image tokens within the same embedding space. The LLMs maintain their original next-token prediction training objective without any changes, with the only variation being that their vocabulary space has been extended from purely language-space to a combined space for both visual and linguistic modalities.
We use the GEMMA-7B [60] model as our primary model to validate its multimodal understanding, image generation capabilities, and performance on text-only tasks after augmenting it with the ability to understand and generate images. To further investigate whether a trade-off exists when accommodating visual generation tasks and text generation tasks within the same LLM space, as well as to explore the scaling performance of different sized models, we conducted same training on LLAMA-3 1B  [14] and GEMMA-2 series [61] at scales of 2B, 9B, and 27B, thereby observing their distinct performances.
3.3 Data Preparation
3.3.1 Continue Pre-training Data
Text-only data. To maintain the language abilities of pre-trained LLMs, we sample text-only data from public datasets for the pre-training stage. We sample 15M text data from DCLM [28], 12M from SlimPajama [55], and 3M code data from Starcoderdata [31], totaling 30M text instances for approximately 60 billion text tokens in the pre-training stage.
Image-text pairs. We use JourneyDB [43] and internal MidJourney-style synthetic data to compile 30M high-quality image data, for a total of 30 billion image tokens. All the data are used to form a hybrid multi-modal data for continue pre-training, resulting in quick acquisition of decent image generation capabilities while retaining language abilities. Furthermore, we reversed 20% of text-to-image data to train in a captioning format to enhance its understanding of images.
3.3.2 Instruction Tuning Data
After the pre-training stage, the model has already attained substantial image generation capabilities and a decent level of image understanding. We use 1M LMSYS [74] as text-only instruction data, coupled with 1M text-to-Image data sampled from high-quality data, and 1.5M multi-modal instruction tuning data introduced in Minigemini [32]. This compiles a 3.5M hybrid instruction tuning data for further refining our model.
3.4 Training Procedure
Data Formulation. During the continue pre-training stage, we define the input format for multi-modal training data as:
@@ 3.4 Training Procedure
Here, [bos] and [eos] are the begin-of-sequence and end-of-sequence tokens defined in the original text tokenizer. We incorporate two additional spatial tokens, [boi] and [eoi], to specifically signify the start and end of image tokens. All candidate images are filtered based on their aspect ratios, retaining those with a ratio smaller than 2, and after center cropping, they are resized to $512\times 512$ pixels. With the image tokenizer, each image has a token length of 1024. During the instruction tuning stage, we construct a series of prompts to guide the model to generate images and randomly select from these during each training iteration.
Training Objective. Since all images are tokenized into discrete image tokens that share the same vocabulary and embedding space as text tokens, the training objective is wholly consistent with LLMs via next-token prediction and utilizes the standard cross-entropy loss. Moreover, we observe that during the continue pre-training stage, models of 7B size and above tend to encounter loss spikes in the early stages of training. It significantly impacts the training performance and convergence speed. To mitigate this issue, we reduce the max grad norm to 0.5 for larger models and use max-z loss [70] to normalize the logits, enhancing the stability of training.
Training Details. In all experiments, we initiate with a learning rate of 2e-5. Except for the scaling experiments that employed a constant learning rate, all other tests utilized a cosine learning schedule. The training batch size is set at 1024, and the max context length is set to 2048. We use the DeepSpeed Zero3 optimization strategy.
@@ 4 Experiments
4.1 Training Details and Evaluation Setting
As Liquid extends the base capabilities of LLMs to generate and understand images, we validate its performance across three aspects: text-guided image generation, visual understanding, and general text-only capabilities.
To evaluate the visual understanding capabilities, we report results on widely-adopted zero-shot image-based benchmarks, which include VQA-v2 [20], GQA [23], TextVQA [54], MME [16].
[…]
SD-XL [44] Diffusion 2000M 9.55
PixArt [6] Diffusion 25M 6.14
Playground v2.5 [27] Diffusion 4.48
Show-o [69] Discrete Diff. 36M 15.18
LWM [39] Autoregressive 17.77
VILA-U (256) [68] Autoregressive 15M 12.81
VILA-U (384) [68] Autoregressive 15M 7.69
Janus [67] Autoregressive - 10.10
Ours Autoregressive 30M 5.47
4.2 Visual Generation Results
@@
In Tab. 1, we report FID on MJHQ-30K to compare the images quality generated by Liquid with other models. It is observable that Liquid not only has a lower FID than all other auto-regressive methods but also surpasses most well-known diffusion models except Playground v2.5 [27]], achieving a very low FID of 5.47. It indicates that LLMs are also capable of generating high-quality images, as shown in Fig. 2, providing proof that the upper limit of LLMs in terms of image aesthetic quality is not inferior to diffusion models.
Due to the LLMs’ capability to output dynamically-length text in the form of next-token prediction, the convenience can be applied to visual generation. We find that by appending instructions about the resolution to the input text prompt, such as "length is: width is:", the model can quickly learn to generate the corresponding code according to the specified number of rows and columns. Fig. 2 demonstrates the generation results at various resolutions, showcasing the flexibility of Liquid.
@@ 4.3 Vision-Language Understanding
To validate the vision-language understanding capabilities of Liquid after instruction tuning, we test our model across various public vision-language benchmarks, as shown in Tab. 3. Compare with the MLLMs with discrete visual token, Liquid outperforms models with stander VQVAE like LWM [39], Chameleon [59], and Show-o [69], and achieves comparable results with VILA-U [68], which have trained a CLIP based multi-codebook VQVAE to improve understanding. However, the performance of MLLMs using discrete visual tokens on visual understanding tasks tends to be lower than mainstream models that employ continuous visual tokens. Most MLLMs with continuous visual token use CLIP features as visual input, it is a significant advantage considering that CLIP is pre-trained on a large-scale image-text pair dataset, leading to a strong alignment between its visual space features and language space features.This alignment substantially aids the LLMs, making it easier for them to understand visual content. In contrast, using image tokens derived directly from VQVAE tokenizer as input means that the corresponding embedding features in the LLM are reinitialized with out any alignment. Without extensive pre-training to align feature spaces, the visual understanding capabilities might be slightly inferior to models using CLIP as visual input. The difference in performance mainly stems from the fact that most VQVAE currently do not align image-text spaces. VILA-U [68] has confirmed that by adding CLIP loss during the VQVAE training to align its visual space, the performance of visual understanding tasks can be improved.
@@
However, we attempt to increase the amount of image-text pairs the model can see during the pre-training phase by training one more epoch on our multi-modal pre-training data, and then the model could achieve better performance on visual understanding tasks. This result indicates that mixed-modality pre-training plays a role similar to CLIP pre-training, aligning text and visual embedding spaces. More pret-raining or more suitable embedding initialization methods could further boost the performance of discrete visual tokens on comprehension tasks.
Method BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c MMLU
Mistral 7B 84.7 83.0 - 81.3 75.3 80.0 55.5 60.1
8$\times$7B - 83.6 - 84.4 77.2 83.1 59.7 70.6
LLAMA2 7B 77.4 78.8 48.3 77.2 69.2 75.2 45.9 45.3
13B 81.7 80.5 50.3 80.7 72.8 77.3 49.4 54.8
34B 83.7 81.9 50.9 83.3 76.7 79.4 54.5 62.6
Chameleon 7B 81.4 79.6 57.0 74.2 70.4 76.1 46.5 52.1
34B 86.0 83.3 63.3 82.7 78.5 84.1 59.7 65.8
Ours 7B 81.0 81.0 46.7 76.1 72.7 75.6 49.0 56.0
4.4 Language Tasks
We evaluate the general text-only capabilities of our pre-trained model against other state-of-the-art large language models and multi-modal language models, following the evaluation protocol of  [17]. As shown in Tab. 4, Liquid outperforms the well-established language model LLAMA2 [64] and the mix-pretrained multi-modal language model Chameleon [59] in most tasks, exhibiting undegraded linguistic capabilities. Compared with Chameleon [59], which is mixed pretrained with an extremely large scale of data, Liquid performs training from existing LLMs that already possess decent language capabilities, maintaining these capabilities without degradation. This result validates the efficiency of our training framework and demonstrates that with this framework, we can extend the visual generation and understanding capabilities to LLMs of any structure and size.
4.5 Scaling Results
Benefiting from our strategy of continuing training from existing LLMs, we can conveniently scale up our model size and observe their performance on different tasks. We construct Liquid models of sizes 1B, 2B, 9B, and 27B using models from the LLAMA3.2 [14] and GEMMA2 [61]. As far as we know, no prior work has explored whether training image and text generation in a single embedding space through a single model would impact their respective performance. To explore this, we conducted three sets of experiments with the same training strategy on the four model sizes: mix image generation and text generation, text generation only, and image generation only. In mix training, we use both 30M text only and 30M image-text data introduced in Sec. 3. For text only and image only training, we only use one of them. All experiments control the batch size to 512, and the learning rate is set to a constant 2e-5. After the training is completed, we evaluate the performance of the three settings and their respective four sizes, totaling 12 models. For language ability, we evaluate two indicators, namely the zero-shot MMLU and the average zero-shot score in 5 language tasks. For visual generation tasks, we evaluate the VQAscore on advanced prompts in GenAI-Bench. As shown in Fig. 1, we find a trade-off phenomenon when both tasks were mixed in the training of smaller models; however, as the model size increased, this trade-off gradually disappeared. This confirms that larger models possess sufficient capacity to handle both visual and language spaces generation concurrently. Furthermore, as the size of the model increases, the quality of the images generated also improves, proving that LLMs have a higher upper limit as multi-modal generators, and the same scaling-law as language models. More scaling results can be found in the supplementary materials.
4.6 Mutual Boost of Visual Generation and Understanding

### v2

@@ Abstract
We present Liquid, an auto-regressive generation paradigm that seamlessly integrates visual comprehension and generation by tokenizing images into discrete codes and learning these code embeddings alongside text tokens within a shared feature space for both vision and language. Unlike previous multimodal large language model (MLLM), Liquid achieves this integration using a single large language model (LLM), eliminating the need for external pretrained visual embeddings such as CLIP. For the first time, Liquid uncovers a scaling law that performance drop unavoidably brought by the unified training of visual and language tasks diminishes as the model size increases. Furthermore, the unified token space enables visual generation and comprehension tasks to mutually enhance each other, effectively removing the typical interference seen in earlier models. We show that existing LLMs can serve as strong foundations for Liquid, saving 100× in training costs while outperforming Chameleon in multimodal capabilities and maintaining language performance comparable to mainstream LLMs like LLAMA2. Liquid also outperforms models like SD v2.1 and SD-XL (FID of 5.47 on MJHQ-30K), excelling in both vision-language and text-only tasks. This work demonstrates that LLMs such as Qwen2.5 and GEMMA2 are powerful multimodal generators, offering a scalable solution for enhancing both vision-language understanding and generation. The code and models will be released at https://github.com/FoundationVision/Liquid.
@@ 1 Introduction
In light of the aforementioned issues, an emerging line of research [15, 50, 12, 72, 63, 57] attempts to employ the VQVAE [66, 15] model as a universal visual tokenizer for MLLMs. Analogous to the role played by the BPE [54] tokenizer in LLMs, VQVAE establishes a bi-directional mapping between raw pixels and discrete codes. This enables the MLLMs to learn visual code embeddings jointly with text tokens, rather than constrained by the feature space of a pretrained visual encoders like CLIP or pretrained visual generators like diffusion models. Moreover, the discrete nature of visual tokens allows uniform modeling of visual and text tokens with the same next-token prediction loss, which seamlessly integrates both modalities. LWM [40] and Chameleon [60] are the pioneers in exploring this approach. However, these methods involve extensive training from scratch, which makes it computationally expensive to disseminate exploration in this form. Follow-up works further introduces bi-directional attention [76], discrete diffusion modeling [71], and residual codes [70] to handle visual tokens. EMU3 [67] has verified that the next prediction paradigm can achieve SOTA results in both understanding and generation tasks by finetuning separate two models, but there is a lack of exploration of the effects of unifying these two tasks within one model. In this paper, we explore the potential of solely employing LLMs as multimodal generators, and conduct a comprehensive empirical study to investigate a series of properties including the scaling laws of visual generation, the impact on original language capabilities, and the interrelation between visual understanding and generation tasks.
[…]
Our experiments reveal several insightful properties: 1) Directly employing LLMs for visual generation exhibits clear scaling laws in both validation loss and image consistency metrics, consistent with those seen in LLMs, regardless of whether the models retain their language capabilities or focus solely on visual generation tasks. 2) When trained with multimodal data such as image-text pairs and pure language data, the language capabilities of the models are lower compared to those trained with only language data. However, this tradeoff diminishes as the model size increases, indicating that larger models have sufficient capacity to handle both tasks seamlessly. This, combined with the scaling law, demonstrates the strong potential of LLMs as unified multimodal generator. 3) The visual understanding and visual generation tasks can mutually benefit each other. We find that increasing the data for either visual generation or visual understanding tasks improves the performance of the other, demonstrating the advantages and natural rationale of this paradigm in joint optimization.
@@
  • Comprehensive experiments about scaling laws of unified multi-modal models, revealing that the trade-off between language and visual tasks diminishes as scale of model increases.
@@
2 Preliminaries
Image Tokenizer. We use the same tokenizer as Chameleon [60], which is a VQGAN [15]. It encodes a 512 × 512 image into 1024 discrete tokens in a codebook of size 8192. We appended these discrete image tokens to the text codebook produced by the BPE [54] tokenizer to expand the vocabulary size of the LLMs, thereby extending its language space to a multi-modal space encompassing both visual and linguistic elements.
Architecture. As illustrated in Fig. 1, Liquid can be constructed based on any existing LLMs. In this paper, we employ the GEMMA-7B [61] as base model of Liquid to validate its multimodal understanding, image generation capabilities, and performance on text-only tasks after augmenting it with the ability to understand and generate images. Then we conduct a comprehensive empirical study on the scaling behavior at scales of 0.5B, 1B, 2B, 7B, 9B, and 32B models from LLAMA-3 [14], GEMMA-2 [62], and Qwen2.5 [24]. We refrain from altering any structures within the LLMs to facilitate continued training directly from pre-trained weights. The only modification is the addition of 8192 new learnable embeddings for discrete image tokens. The LLMs maintain their original next-token prediction training objective with the standard cross-entropy loss.
Data Preparation. To maintain the language abilities of pre-trained LLMs, we sample text-only data from public datasets for the pre-training stage. We sample 15M text data from DCLM [28], 12M from SlimPajama [56], and 3M code data from Starcoderdata [31], totaling 30M text instances for approximately 60 billion text tokens. For image-text pairs, we use JourneyDB [44] and internal MidJourney-style synthetic data to compile 30M high-quality image data, for a total of 30 billion image tokens. All the data are used to form a hybrid multi-modal data for continue pre-training, resulting in quick acquisition of decent image generation capabilities while retaining language abilities. Furthermore, we reversed 20% of text-to-image data to train LLMs in a captioning task to enhance their visual understanding abilities.
Training Procedure. We use a total of 60M data to continue pretraining our models. For all image-text data, we define the input format for multi-modal training data as:
@@ 2 Preliminaries
Here, [bos] and [eos] are the begin-of-sequence and end-of-sequence tokens defined in the original text tokenizer. We incorporate two additional spatial tokens, [boi] and [eoi] to signify the start and end of image tokens. In the scaling experiments, for each model size, we independently use 30M text-only data, 30M text-to-image data, and a mixed dataset of 60M data to train three different versions of the models. We then evaluate their performance on a series of tasks.
@@ 3 Experiments
3.1 How Well Do LLMs Perform in Visual Generation?
In this section, we comprehensively evaluate the visual generation capabilities of Liquid-7B model that have been continuously pretrained with a combined dataset of 60M mixed multimodal data. Subsequently, we evaluate mixed pre-trained models across six different sizes, ranging from 0.5B to 32B, the objective is to explore the scaling behavior of LLMs in visual generation tasks, while simultaneously preserving their linguistic capabilities.
@@ 3.1 How Well Do LLMs Perform in Visual Generation?
3.1.1 Quantitative Results on Visual Generation Benchmarks
[…]
SD-XL [45] Diffusion 2000M 9.55
PixArt [6] Diffusion 25M 6.14
Playground v2.5 [27] Diffusion 4.48
Show-o [71] Discrete Diff. 36M 15.18
LWM [40] Autoregressive 17.77
VILA-U (256) [70] Autoregressive 15M 12.81
VILA-U (384) [70] Autoregressive 15M 7.69
Janus [69] Autoregressive - 10.10
Liquid-7B Autoregressive 30M 5.47
@@ 3.1.1 Quantitative Results on Visual Generation Benchmarks
In Tab. 2, we report FID on MJHQ-30K to compare the images quality generated by Liquid with other models. It is observable that Liquid not only has a lower FID than all other auto-regressive methods but also surpasses most well-known diffusion models except Playground v2.5 [27], achieving a very low FID of 5.47. It indicates that LLMs are also capable of generating high-quality images, as shown in Fig. 2, providing proof that the upper limit of LLMs in terms of image aesthetic quality is not inferior to diffusion models. Further more, due to the capability of LLM to generate dynamically-length content in the form of next-token prediction, the convenience can be applied to visual generation. We find that by appending instructions about the resolution to the input text prompt, such as "length is: width is:", the model can quickly learn to generate the corresponding code according to the specified number of rows and columns. Fig. 2 demonstrates the generation results at various resolutions, showcasing the flexibility of Liquid.
Figure 3: Scaling behavior of LLMs in multimodal generation across different sizes. We explore the visual generation performance of LLMs ranging from 0.5B to 32B in size after undergoing mixed training with language data and text-to-image data. The validation loss smoothly decreases as both the model size and training iterations increase, while token accuracy and VQA Score consistently rise. The VQA Score consistently increases as the validation loss decreases, indicating a strong correlation between them.Uder the same training FLOPs, smaller models reach lower validation losses more quickly, but larger models ultimately achieve higher evaluation metrics.
3.1.2 Scaling Results on Visual Generation
We explore the visual generation performance of LLMs ranging from 0.5B to 32B in size after mixed training with language data and text-to-image data. As shown in Fig. 3, with the increase in model size and training iterations, the validation loss smoothly decreases, while token accuracy and VQA Score consistently increase. Moreover, the VQA score continues to rise as the validation loss decreases, indicating a strong correlation between validation loss and holistic image evaluation metrics. Under the same training FLOPs, smaller models achieve lower validation losses and higher VQA Scores more quickly, but larger models ultimately reach higher evaluation metrics. This may be due to smaller models being able to iterate through more training steps rapidly, allowing for swift changes in model capabilities. Although smaller models can quickly acquire visual generation abilities, their upper limits are lower, making it difficult to achieve high-quality visual generation results. As shown in Fig. 4, larger models eventually obtain more robust visual generation outcomes.
Figure 4: Scaling model size and training compute improves visual fidelity and soundness. Zoom in for a better view. Samples are drawn from Liquid models of 4 different sizes (0.5B, 1B, 2B, 9B) and 3 different training steps (5K, 15K, 40K).
3.2 Does Visual Generation Compromise Language Ability?
In this section, we evaluate the impact on the inherent linguistic capabilities of LLMs after they acquire visual generation abilities. First, we compare the continuously pretrained GEMMA-7B base model with mainstream language models to determine if it has retained robust linguistic capabilities. Subsequently, we evaluate LLMs of different sizes that have been trained either exclusively with 30M language data or with a mix of 60M multimodal data. Since both training processes use the same language data, this allows for a fair comparison of how different sizes of LLMs are affected in terms of their original linguistic capabilities after acquiring visual generation abilities.
3.2.1 Comparison with Mainstream LLMs
We evaluate the general text-only capabilities of our mixed multimodal pretrained model against other state-of-the-art large language models and multi-modal language models, following the evaluation protocol of  [17]. As shown in Tab. 3, Liquid outperforms the well-established language model LLAMA2 [65] and the mix-pretrained multi-modal language model Chameleon [60] in most tasks, exhibiting undegraded linguistic capabilities. Compared with Chameleon [60], which is mixed pretrained with an extremely large scale of data, Liquid performs training from existing LLMs that already possess decent language capabilities, maintaining these capabilities without degradation. This result validates the efficiency of our training framework and demonstrates that with this framework, we can extend the visual generation and understanding capabilities to LLMs of any structure and size.
Method BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c MMLU
Mistral 7B 84.7 83.0 - 81.3 75.3 80.0 55.5 60.1
8$\times$7B - 83.6 - 84.4 77.2 83.1 59.7 70.6
LLAMA2 7B 77.4 78.8 48.3 77.2 69.2 75.2 45.9 45.3
13B 81.7 80.5 50.3 80.7 72.8 77.3 49.4 54.8
34B 83.7 81.9 50.9 83.3 76.7 79.4 54.5 62.6
Chameleon 7B 81.4 79.6 57.0 74.2 70.4 76.1 46.5 52.1
34B 86.0 83.3 63.3 82.7 78.5 84.1 59.7 65.8
Liquid-7B 7B 81.0 81.0 46.7 76.1 72.7 75.6 49.0 56.0
Figure 5: The performance impact of multi-modal mixed training versus text-only training on language tasks. Multi-modal mixed training does impact language performance when the model size is small. However, this degradation gradually disappears as the model size increases.
3.2.2 Scaling Results on Language Tasks
To investigate whether visual generation capabilities impact linguistic abilities, we compare the performance of models trained with 30M language data and those trained with a mix of 60M multimodal data on language tasks across different sizes. For language ability, we report the average average scores on 5 language tasks, including HellaSwag, WinoGrande, ARC-Easy, ARC-Challenge, and BoolQ. As shown in Fig. 5, we find a trade-off phenomenon when both tasks were mixed in the training of smaller models; however, as the model size increased, this trade-off gradually disappeared. This confirms that larger models possess sufficient capacity to handle both visual and language spaces generation concurrently.
3.3 Does Language Ability Limit Visual Generation Performance?
In language tasks, we observe a trade-off in performance of multimodal LLMs, raising the question of whether this trade-off also exists in visual generation tasks. To address this question, we compare LLMs of different sizes that have been trained exclusively with text-to-image data and with a mix of multimodal data. Both training utilize the same visual generation data, with the only difference being whether the models retain their linguistic capabilities.
Figure 6: Scaling results between T2I only training and Mix training. Mix training results in higher validation loss on visual generation tasks for models of every size. However, its impact on the VQA Score diminishes as the size of the model increases.
As illustrated in Fig. 6, mixed training leads to higher validation loss on visual generation tasks for models of every size. However, the detrimental effect on the VQA Score decreases as the model size increases. This reduction is partly due to the increased capacity of larger models and may also be attributed to the evaluation metrics approaching saturation. Nevertheless, it can be argued that as the size of the model increases, the trade-off between visual generation and language tasks gradually diminishes or even disappears. This encourages leveraging the scaling capabilities of LLMs into multimodal models.
3.4 How Does Liquid Perform on Visual Understanding Tasks?
After the pre-training stage, the model has already attained substantial image generation capabilities and a decent level of image understanding. To evaluate the visual understanding capabilities, we use 1M LMSYS [75] as text-only instruction data, coupled with 1M text-to-Image data sampled from high-quality data, and 1.5M multi-modal instruction tuning data introduced in Minigemini [33]. This compiles a 3.5M hybrid instruction tuning data for further refining our model. We report results on widely-adopted zero-shot image-based benchmarks, which include VQA-v2 [20], GQA [23], TextVQA [55], POPE [32], and MME [16].
@@ 3.4 How Does Liquid Perform on Visual Understanding Tasks?
As shown in Tab. 4, compare with the MLLMs with discrete visual token, Liquid outperforms models with stander VQVAE like LWM [40], Chameleon [60], and Show-o [71], and achieves comparable results with VILA-U [70], which have trained a CLIP based multi-codebook VQVAE to improve understanding. However, the performance of MLLMs using discrete visual tokens on visual understanding tasks tends to be lower than mainstream models that employ continuous visual tokens. Most MLLMs with continuous visual token use CLIP features as visual input, it is a significant advantage considering that CLIP is pre-trained on a large-scale image-text pair dataset, leading to a strong alignment between its visual space features and language space features.This alignment substantially aids the LLMs, making it easier for them to understand visual content. In contrast, using image tokens derived directly from VQVAE tokenizer as input means that the corresponding embedding features in the LLM are reinitialized with out any alignment. Without extensive pre-training to align feature spaces, the visual understanding capabilities might be slightly inferior to models using CLIP as visual input. The difference in performance mainly stems from the fact that most VQVAE currently do not align image-text spaces. VILA-U [70] has confirmed that by adding CLIP loss during the VQVAE training to align its visual space, the performance of visual understanding tasks can be improved.
@@
However, we attempt to increase the amount of image-text pairs the model can see during the pre-training phase by training one more epoch on our multi-modal pre-training data, and then we were surprised to discover that the model could achieve better performance on visual understanding tasks. This result indicates that mixed-modality pre-training plays a role similar to CLIP pre-training, aligning text and visual embedding spaces. More pret-raining or more suitable embedding initialization methods could further boost the performance of discrete visual tokens on visual understanding tasks.
3.5 Will Understanding and Generation Tasks Mutually Improve Each Other?
As observed in above experiments on visual understanding tasks, extended pretraining enhances comprehension abilities. Notably, during the pretraining process, 80% of the data was trained in a text-to-image format. This observation naturally leads us to question whether increased training with text-to-image data could improve performance on understanding tasks. Furthermore, we explore whether there is a synergistic relationship between these two tasks.